How long will it take a mother, running at 4 feet per second, to catch up with her toddler, running down the sidewalk at 2 feet per second, if the child had a 5 second head start
Answer by lwsshak3(11628)

How long will it take a mother, running at 4 feet per second, to catch up with her toddler, running down the sidewalk at 2 feet per second, if the child had a 5 second head start
***
travel time*speed=distance
5*2=10 ft=distance mother must catch up with her toddler
4-2=2 ft/s=speed at which mother is catching up to her toddler
10/2=5sec
How long will it take a mother, running at 4 feet per second, to catch up with her toddler? 5 sec